OpenStudy (anonymous):

derivative of (2x + 1)^3

OpenStudy (anonymous):

is the 2x considered an internal function, if so, I would multiply 3(2x + 1)2 by 2?

OpenStudy (solomonzelman):

Example: \(\large\color{black}{ y= (4x + 1)^4 }\) the derivative would be: \(\large\color{black}{ y'= 4(4x + 1)^{4-1} \times (4x+1)' }\) \(\large\color{black}{ y'= 4(4x + 1)^{3} \times 4}\) \(\large\color{black}{ y'= 16(4x + 1)^{3}}\)

OpenStudy (anonymous):

@cleetus779 You are correct

OpenStudy (uri):

Yes, that is correct. Good job!

OpenStudy (anonymous):

Here is my result:\[ d(2x + 1)^3 = du^3 = 3u^2du = 3(2x+1)^2d(2x+1) = 6(2x+1)^2dx\\ \implies \frac{d(2x+1)^3}{dx} = 6(2x+1)^2 \]

OpenStudy (solomonzelman):

to me it looks even a bit more confusing using a "u"... but everyone has a different taste.

OpenStudy (anonymous):

many thanks to all, the problem actually asks for the derivative of tan^2((2x + 1)^3), can you help me continue?

OpenStudy (solomonzelman):

Ohh, sure: Can you tell me what is the derivative of tan(x) ?

OpenStudy (anonymous):

I get the 'u' stuff, it actuallyhelps to visualize it differently, so I appreciate that

OpenStudy (solomonzelman):

what is the derivative of tan(x) ?

OpenStudy (anonymous):

sec^2(x)

OpenStudy (solomonzelman):

yes.

OpenStudy (anonymous):

now product rule?

OpenStudy (solomonzelman):

So what you are getting is. \(\large\color{black}{ \tan[~(2x + 1)^3 ~] }\) the derivative would be: \(\large\color{black}{ \sec^2~[~(2x + 1)^3 ~] \times 3(2x + 1)^2 \times 2 }\)

OpenStudy (solomonzelman):

see how there is there multiple chains going on?

OpenStudy (anonymous):

processing

OpenStudy (solomonzelman):

(you are using chain rule twice)

OpenStudy (anonymous):

going to work out solution, will post what i get, could you do the same?

OpenStudy (solomonzelman):

I already have the solution

OpenStudy (solomonzelman):

anyway, work it out, and we will see if anything should be corrected....